# Signature Kernel Scoring Rule as Spatio-Temporal Diagnostic for Probabilistic Forecasting

### v1

@@ A.4 RQE
RQE measures the overall forecast tendency across all extreme quantiles, with RQE << 0 implying a trend of underestimation of extremes, while RQE >> 0 implies an overestimation of extremes compared to the true distribution. However, this value is limited by not assessing whether extreme values are predicted correctly, as it is insensitive to correct alignment with observations. As a result, this value needs to be interpreted along with an assessment of forecast accuracy [Bi et al., 2023].
@@ A.5 Calibration
We consider the assessment of calibration for probabilistic forecasts to be a relevant measure for inclusion in evaluation. We use the definition from Radev et .al of calibration error as the difference between credible intervals in the forecast and observation distributions [Radev et al., 2020]. For each variable, we consider 100 credibility levels $\alpha$ spaced evenly between 0.01 and 1. We compute the $\alpha/2$ and $1-\alpha/2$ quantiles from our forecast samples $f^{v}$, $q_{a/2}^{v}$ and $q_{1-a/2}^{v}$ respectively. For each credible level we determine the proportion of true samples which lie within the corresponding credible interval, known as the empirical coverage or proportion of inliers:
@@
Appendix B Evaluation Implementation
@@
There are two different implementation approaches taken for either evaluation or model training, both relating to latitude weighting and stability of the signature kernel. Training implementation and the discussion of generative model architecture and the prequential is found in the main body in Section 4.
@@
Here we address evaluation, where latitude weighting is critical for appropriate model assessment. Including spatial correlations within the signature kernel score while latitudinally weighting requires additional considerations. Latitude weighting is necessary to prevent the poles from having a disproportionate contribution to the model compared to the equator. However, due to the non-linear transformation of the signature, and the implicit mapping via the kernel, different latitude regions can’t be appropriately pre-weighted before applying the signature kernel. Instead, from each initialisation time, multiple paths are considered, each corresponding to a latitude slice with dimensionality equal to the number of longitude positions at every time step. From each path, the relevant values to calculate the score or distance are determined between the forecasted data X and observations Y: $K^{Sig}(X,Y),K^{Sig}(X,X),K^{Sig}(Y,Y)$. A weighted average is then able to be conducted on these values associated with each latitude slice, before calculating the score or distance. This requires a revisit to the signature kernel definitions. In the simpler deterministic case, we consider a forecast $F_{i,j,t,k}$ with longitude, latitude $i,j$, with respective number of each $I,J$, the set of initialisation times {tt}, and future predictions $1:k$ for each initialisation time, along with observations $o_{i,j,t}$ , with likewise $i,j$ and values at {tt} + $1:k$. We consider the latitude weighted signature distance and score for a path length of $k\geq 2$ over the set of initialisation times {tt}, as follows:
@@
Variables measured at different atmospheric pressures, commonly 500 hPa and 850 hPa, which correspond to differing three-dimensional areas, are treated independently. The summation over longitude slices removes the spatial correlations between different latitudes from being captured by the signature kernel during evaluation. However, without latitude separation, regions are not compared appropriately. Hemisphere and tropical separations which are required for scorecard analysis are created via assigning the appropriate latitude slices.
Appendix C Numerical Instability
The largest issue with implementation is the instability of computation, with different scales and static kernels causing issues with Inf/Nan values or scores converging to 1. In multiple test instances, the signature kernel score appears to be calculating correctly for small prediction lags, but then displays abnormal converging behaviour. Theoretical and experimental results were limited in application to this context. The generalised data augmentation method proposed by James Morril et al. [Morrill et al., 2021] demonstrated results of different approaches varying significantly by dimensionality of the path, with eight being the largest dimension tested. Moreover, results were only conducted for path length $l>>d$, the dimension of the data. In this implementation, the dimensionality of the path is 2048 in evaluation and 256 in training.
This type of path is not covered by existing literature. Nonetheless, extending these results proved successful. Standard normalising using the mean and standard deviation of only the observation data is conducted to allow comparisons between models with equal scaling. A further scaling constant of $\frac{1}{\sqrt{I\cdot J}}$ is applied, which is one over the square root of the longitude and latitude dimensions. Despite no theoretical backing, and with pre-processing rescaling being primarily considered for the truncated signature instead [Morrill et al., 2021], this addition solved detectable instability issues for ensemble models and across increasing resolution, verified up to the standard 240 by 121 resolution. The RBF kernel is considerably more robust to numerical instability issues in both evaluation and training than other kernels tested. The linear identity kernel was initially desired for handling high dimensional time series without complex feature mapping, to simplify the interpretation of the kernel, but numerical instability issues remain in training.
Appendix D WeatherBench data
@@
WeatherBench 2 is an open-source evaluation frame-work which contains the current state-of-the-art models along with training, ground truth, and baseline data [Rasp et al., 2020]. All models compared in this work originate from the WeatherBench 2 system. There are two distinct weather datasets considered by WeatherBench models, either Analysis or ERA5 reanalysis data. Both are created via the ECMWF’s 4D-VAR data assimilation as best guesses of the true state of the atmosphere. However, Analysis is created in real time as the operational best predictor of the atmosphere. ERA5 reanalysis is a dataset created in retrospect, combining observations and ECWMF’s HRES model predictions over twelve hour windows. In general, ERA5 data can be assumed to be closer to the truth than Analysis data, but operational models benefit from training with Analysis data. In either case, each model should be evaluated against the respective data set it was trained on, with significant differences only present for early lead times [WeatherBench, 2023]. For this work, we have only considered ERA5 reanalysis data.
@@
ERA5 data exists for each hour since 1940, however all current models only train on data since 1979. This year marks a significant jump in accurate weather observation, particularly for the Southern Hemisphere, relative to previous years, as global satellite weather monitoring was introduced [Uppala, 2007]. Additionally, most models consider a down-sampled version, with data every six hours at 00:00, 06:00, 12:00, and 18:00 UTC. Forecasts are initialised either on this schedule, or on a twelve hour cycle 00:00, 12:00. Predictions are then given every six or twelve hours. Forecasts are split between providing a zero-time initialisation or not, i.e. giving a prediction for its initialisation time. When comparing models with different zero-time forecasting approaches, this forecast was removed.
@@
ERA5 reanalysis data between 1979 and roughly three months behind the present exists in completion without missing data at a 0.25 latitude-longitude coordinate resolution. It can be downloaded in its entirety from the Copernicus Climate Data store [Hersbach et al., 2023]. Due to the extensive work in creating this dataset, there is no need to address any data integrity issues. Predictions from WeatherBench models can be downloaded from the WeatherBench platform. Certain models do have results missing for certain variables or resolutions, along with missing predictions. Variables or resolutions with missing data were not considered. The signature kernel method can handle missing data via the time augmentation, and so no further action was taken.

### v2

@@ A.4 RQE
RQE measures the overall forecast tendency across all extreme quantiles, with RQE << 0 implying a trend of underestimation of extremes, while RQE >> 0 implies an overestimation of extremes compared to the true distribution. However, this value is limited by not assessing whether extreme values are predicted correctly, as it is insensitive to correct alignment with observations. As a result, this value needs to be interpreted along with an assessment of forecast accuracy (Bi et al., 2023).
@@ A.5 Calibration
We consider the assessment of calibration for probabilistic forecasts to be a relevant measure for inclusion in evaluation. We use the definition from Radev et .al of calibration error as the difference between credible intervals in the forecast and observation distributions (Radev et al., 2020). For each variable, we consider 100 credibility levels $\alpha$ spaced evenly between 0.01 and 1. We compute the $\alpha/2$ and $1-\alpha/2$ quantiles from our forecast samples $f^{v}$, $q_{a/2}^{v}$ and $q_{1-a/2}^{v}$ respectively. For each credible level we determine the proportion of true samples which lie within the corresponding credible interval, known as the empirical coverage or proportion of inliers:
@@
Appendix B Geometric Signature Interpreations
$S(X)^{i,i}_{a,b}=\frac{(X^{i}_{b}-X^{i}_{a})^{2}}{2}=\frac{(S(X)^{i}_{a,b})^{2}}{2}$ (5)
However, non-diagonal terms, referred to as cross terms, with $i_{1}\neq i_{2}$ for second order terms are more complicated, relating to the Lévy area of the two-dimensional path. The Levy area is the signed area between a path and a chord created between its endpoints, represented in Figure B.1 (Chevyrev and Kormilitzin, 2025) with areas $A_{-}$ and $A_{+}$, denoted by:
[…]
Figure B.1: Representation of the signed Lévy area of a two-dimensional path. $A_{+}$ and $A_{-}$ denote the positive and negative respective areas between the path (red) and chord (blue). $S^{1,2}$ and $S^{2,1}$ denote the second order terms equal to the area between the path and their perpendicular enclosure (black).
We can extend the results found in Equations 5 and 6 using the shuffle product identity (Ree, 1958). This identity shows that we can relate the product of any two signature terms $S(X)_{a,b}^{i_{1},\ldots,i_{k}}$ and $S(X)_{a,b}^{j_{1},\ldots,j_{m}}$ as the sum of $S(X)^{K}_{a,b}$ terms, which depend only on indexes ($i_{1},\ldots,i_{k}$) and ($j_{1},\ldots,j_{m})$. This collection of new indexes, $K$, is obtained by permuting both index sets together while keeping their respective ordering, referred to as shuffling and denoted by . Given a permutation $\sigma$ of the set $i_{1},\ldots,i_{k},j_{1},\ldots,j_{m}$}, $\sigma$ is a $(k,m)$-shuffle if $\sigma^{-1}(i_{1})<\ldots<\sigma^{-1}(i_{k})$ and $\sigma^{-1}(j_{1})<\ldots<\sigma^{-1}(j_{m})$. We can therefore define the shuffle product identity:
$S(X)^{i_{1},...,i_{k}}_{a,b}\cdot S(X)^{j_{1},...,j_{m}}_{a,b}=\sum_{K\in\{i_{1},...,i_{k}\}\shuffle\{j_{1},...,j_{m}\}}S(X)^{K}_{a,b}$
Appendix C Extension to Discrete Weather Data
Section 2.2 considered signatures defined for smooth paths. However, in our context, we would like to view our discrete time forecasts and observations as paths. Additionally, with handling real weather data, it is necessary to consider how irregularly sampled or partially observed data affects the result. This requires extending our definitions of the signature from paths to data streams consisting of discrete points. We consider a $d$-dimensional data stream, $X\in\mathbb{R}^{d}$ for $N$ points with an ordered set of times, $t_{0}<t_{1}<\dots<t_{N}$, as follows:
$X=\{\left(X^{1}_{t_{i}},\dots,X^{d}_{t_{N}}\right)\}^{N}_{i=0}$
$\hat{X}_{t}=X_{t_{i}}+\frac{t-t_{i}}{t_{i+1}-t_{i}}(X_{t_{i+1}}-X_{t_{i}})$
Each linear interpolation is piecewise infinitely continuously differentiable $C^{\infty}$ and satisfies the requirements of a smooth path. The overall path is the result of concatenating each interpolated section, resulting in a continuous $C^{0}$ path for any $t\in\left[t_{0},t_{N}\right]$ (Chevyrev and Kormilitzin, 2025). We now revisit the calculation of the signature by considering signatures of each interpolated linear path on their respective interval $t_{i}$, $t_{i+1}$]. Letting $x$ be the straight line defined on $t_{i}$, $t_{i+1}$], we define the signature:
$S(X)_{t_{i}:t_{i+1}}^{j_{1},\ldots,j_{k}}=\frac{1}{k!}\prod^{k}_{l=1}\left(x_{j_{l}}(t_{i+1})-x_{j_{l}}(t_{i})\right)$
$S(X)_{t_{a}:t_{c}}^{j_{1},\ldots,j_{k}}=\sum_{l=0}^{k}S(X)_{t_{a}:t_{b}}^{j_{1},j_{2},\ldots,j_{l-1}}\cdot S(X)_{t_{b}:t_{c}}^{j_{l},j_{l+1},\ldots,j_{k}}$
$S(X)^{1}_{t_{a},t_{c}}=S(X)^{1}_{t_{a},t_{b}}+S(X)^{1}_{t_{b},t_{c}}$
$S(X)^{4}_{t_{a},t_{c}}=S(X)^{4}_{t_{a},t_{b}}+S(X)^{3}_{t_{a},t_{b}}\otimes\ldots$
Now the signature is defined for the region $t_{a}$, $t_{c}$], and can be extended to further points $t_{d}$ with step by step concatenation following the same method. We now have a method to consider weather data and our associated forecasts as paths which can be appropriately transformed into a signature, capturing the dependencies of our chronological sequence across the latitude-longitude grid.
Appendix D Further Signature Augmentations
As the signature extracts information about enclosed areas with the path, this formation directly aligns with this ability by causing loops to appear from the data. This representation benefits the capturing of certain statistics, particularly the quadratic variation, which can be represented as signature terms of the lead-lagged path (Chevyrev and Kormilitzin, 2025):
[…]
An additional choice of windowing operation can be taken. The global window is considered naturally, but changing window size via sliding, expanding or hierarchical dyadic windows changes how data is read to capture information at different scales. There is evidence that the hierarchical dyadic windowing outperforms the global window (Morrill et al., 2021). However, during implementation we observed that large path lengths posed challenges for any strategy of weather forecasting. Over the smaller path lengths that were considered, the global window was deemed suitable. Further investigation into this hyperparameter choice would be a valid extension.
$f(\mathbf{x})-L(S(\mathbf{x}))||\leq\epsilon$
The standard signature provides a natural basis for functions of the time series. In practice, there is no consensus on which version is most suited for a machine learning task. Additionally, as shown with the kernel trick, we do not need to be concerned with lowering the feature dimension of the signature, and so we use the standard signature.
Appendix E Strict Propriety
Proof of Theorem 1.
The score in Equation 1 takes the standard form of a kernel score, which is strictly proper on a compact set of paths $\mathcal{X}$ if and only if $K^{Sig}$ is a characteristic kernel on that space (Steinwart and Ziegel, 2021, Theorem 1.1). We therefore need to verify (i) the compactness of the path space, and that (ii) the signature kernel is characteristic.
(i) Compactness of the path space.
In the discrete forecasting setting, paths are piecewise linear with fixed knots at times $0=t_{0}<\cdots<t_{N}=T$ and values in a compact spatial domain $K\subset\mathbb{R}^{d}$ (closed and bounded via physical properties of weather variables), so the path space is:
$\mathcal{X}=\{x\in C([0,T],\mathbb{R}^{d}):x\text{ is linear on each }[t_{k},t_{k+1}],\;x(t_{k})\in K\;\forall\,k\}.$
We verify the hypotheses of the Arzelà–Ascoli theorem (Rudin, 1976, Thm. 7.25).
  1. 1.
    Uniform boundedness is immediate since every path takes values in $K$.
  2. 2.
    Equicontinuity follows because each path is Lipschitz with the shared constant $L=\operatorname{diam}(K)/\Delta t_{\min}$, where $\Delta t_{\min}=\min_{k}(t_{k+1}-t_{k})$: on every linear segment $\lvert x(t)-x(s)\rvert\leq L\lvert t-s\rvert$, and the bound extends globally over segments by the triangle inequality.
  3. 3.
    Closedness holds because uniform limits of piecewise-linear paths with knots at $t_{0},\dots,t_{N}$ and values in the closed set $K$ retain both properties.
Thus $\mathcal{X}$ is compact in the uniform topology.
(ii) Characteristicness of the signature kernel.
The signature kernel, induced by a characteristic static kernel (such as the radial basis function kernel), is a characteristic kernel on compact sets of bounded-variation paths provided the paths are distinct (Issa et al., 2023, Props. 3.1, 3.3). Bounded variation holds for every $x\in\mathcal{X}$, since each path is Lipschitz. Path uniqueness is guaranteed by the applied basepoint and time signature augmentations (Section 2.3), which removes spatial and temporal invariances.
Conclusion.
Conditions (i) and (ii) together satisfy the hypotheses of Steinwart and Ziegel (2021, Theorem 1.1), so the signature kernel score (1) is strictly proper, exhibiting the desired fundamental property of a scoring rule, ensuring truthful predictions. ∎
Appendix F Evaluation Implementation
@@
There are two different implementation approaches taken for either evaluation or model training, both relating to latitude weighting and stability of the signature kernel. Training implementation and the discussion of generative model architecture and the prequential are found in the main body in Section 4.
@@
Here we address evaluation, where latitude weighting is critical for appropriate model assessment. Including spatial correlations within the signature kernel score while latitudinally weighting requires additional considerations. Latitude weighting is necessary to prevent the poles from having a disproportionate contribution to the model compared to the equator. However, due to the non-linear transformation of the signature and the implicit mapping via the kernel, different latitude regions can’t be appropriately pre-weighted before applying the signature kernel. Instead, from each initialisation time, multiple paths are considered, each corresponding to a latitude slice with dimensionality equal to the number of longitude positions at every time step. From each path, the relevant values to calculate the score or distance are determined between the forecasted data X and observations Y: $K^{Sig}(X,Y),K^{Sig}(X,X),K^{Sig}(Y,Y)$. A weighted average is then able to be conducted on these values associated with each latitude slice, before calculating the score or distance. This requires a revisit to the signature kernel definitions. In the simpler deterministic case, we consider a forecast $F_{i,j,t,k}$ with longitude, latitude $i,j$, with respective number of each $I,J$, the set of initialisation times {tt}, and future predictions $1:k$ for each initialisation time, along with observations $o_{i,j,t}$, with likewise $i,j$ and values at {tt} + $1:k$. We consider the latitude weighted signature distance and score for a path length of $k\geq 2$ over the set of initialisation times {tt}, as follows:
@@
Variables measured at different atmospheric pressures, commonly 500 hPa and 850 hPa, which correspond to differing three-dimensional areas, are treated independently. The summation over longitude slices removes the spatial correlations between different latitudes from being captured by the signature kernel during evaluation. However, without latitude separation, regions are not compared appropriately. Hemisphere and tropical separations, which are required for scorecard analysis, are created by assigning the appropriate latitude slices.
Appendix G Discriminative Power
Discriminative power is the ability of a score to distinguish between different forecast and observation paths. High discriminative ability means the score gives significantly different values to forecasts of different quality. Highly discriminative scores prevent ambiguous results when comparing models and also greatly affect the sensitivity of the loss landscape during training (Pinson and Tastu, 2013).
To investigate discriminative power, we implement a stochastic vegetation and fire spread model. Given a two-dimensional grid, we define the vegetation and heat at time $t$ in the coordinate $(i,j)$ as $V_{i,j}^{t}$ and $H_{i,j}^{t}$ respectively. Every time step has the following three processes: vegetation growth, heat spread, and vegetation consumption, using the rates for vegetation growth,$a$, ignition sensitivity, $b$, combustion, $c$, fire day, $d$, noise scale, $s$, and time step $\Delta t$. We introduce stochasticity into the model with $\epsilon$, $\zeta$ $\sim$ $\mathcal{N}(0,s^{2})$.
$V^{t^{\prime}}_{i,j}=V^{t}_{i,j}+a(1-V^{t}_{i,j})\Delta t+\epsilon^{t}_{i,j}$
$H^{t+1}_{i,j}=H^{t}_{i,j}+(b\cdot\bar{H}^{t}_{i,j}\cdot V^{t^{\prime}}_{i,j}-dH^{t}_{i,j})\Delta t+\zeta^{t}_{i,j}$
$V^{t+1}_{i,j}=V^{t^{\prime}}_{i,j}-cH^{t+1}_{i,j}\Delta t$
$\bar{H}^{t}_{i,j}$ is the average value of distance-weighted adjacent and diagonal squares. As a result, we have a simple model, which we can easily simulate over a large grid, with temporally and spatially complex relationships. Vegetation can be thought of as primarily temporally related, while heat is both spatially and temporally correlated. This model was chosen for low time complexity generation to enable scores over large numbers of ensembles, along with smooth path extensions given any initial input data.
To investigate discriminative power, we consider the following experiments on the particular variables of vegetation growth and fire spread intensity. We consider a dense spectrum of changes to each parameter, and for each shifted value, we create the paths of 50 ensemble members, along with 100 simulations of true path trajectories with an unaltered true parameter. As a result, we can estimate the expected score for each ensemble set associated with a differently valued parameter. For a given parameter, ensemble members result in moderately varied path trajectories via the stochasticity in each time step. We assess the relative discriminative power of the signature kernel as compared to the energy score, average MSE, and the variogram score.
We investigate these scores on changes to vegetation growth in Figure G.2. The x-axis represents the change to the shifted parameter of the particular ensemble set. Therefore, the minimum score is expected at zero when both the ensembles and observation paths have aligned parameters. We find that each score captures this variation similarly, with nearly identical curvature up to differences in scale.
We consider these scores on changes to fire spread intensity in Figure G.3. On this parameter, which is both spatially and temporally correlated, we see the signature kernel score displaying a much more drastic spike in score compared to the alternatives. We additionally see greater increases in score with positive changes to fire spread intensity compared to flattening with negative changes. This is intuitive, as the fire dies out for a range of lower fire spread values, causing similar paths, while the opposite is true for fire propagation. The asymmetric profile of the signature kernel score reflects the physical structure of the model. However, despite high sensitivity in the positive region, avoiding plateaus is particularly relevant for learning, i.e. training a model, as our predicted forecasts need to be directed to the optimum. Despite sharp discriminative ability in comparing similar forecasts near the optimum, i.e. the case of diagnostic evaluation between two well trained models, the plateau regions could result in slower, less optimal learning (Goodfellow et al., 2016). On the other hand, while the MSE and the variogram score display smoother curves, they are relatively flat in the region of the optimum, which makes them less sensitive as a diagnostic tool (Gneiting and Raftery, 2007).
In Figure G.4, we compare the discriminative behaviour of the signature kernel score across three static kernels, the RBF, the linear, and the polynomial kernel $d=2$. Both the linear kernel and the polynomial kernel explode in value, with the polynomial kernel in particular displaying highly unstable behaviour. These suffer from numerical stability with unbounded growth and require significant attention to scaling. While asymmetrical, the RBF kernel score does not explode, with kernel values being bounded by $(0,1]$ by construction. The exponential decay of the RBF kernel makes the behaviour observed qualitatively different from the overflow of the linear and polynomial kernels, with this kernel being the most robust to scale, and most suitable for training.
Figure G.2: Comparison of four scores across changes to the vegetation growth parameter. The vertical dotted line at zero represents the true parameter value, with negative and positive deviations to the left and right, respectively.
Figure G.3: Comparison of four scores across changes to the fire spread parameter. The vertical dotted line at zero represents the true parameter value, with negative and positive deviations to the left and right, respectively. The signature kernel score displays higher discriminative power than the other three.
Figure G.4: Comparison of three static kernels across changes to the fire spread parameter. The vertical dotted line at zero represents the true parameter value, with negative and positive deviations to the left and right, respectively.
Appendix H WeatherBench data
@@
WeatherBench 2 is an open-source evaluation frame-work which contains the current state-of-the-art models along with training, ground truth, and baseline data (Rasp et al., 2020). All models compared in this work originate from the WeatherBench 2 system. There are two distinct weather datasets considered by WeatherBench models, either Analysis or ERA5 reanalysis data. Both are created via the ECMWF’s 4D-VAR data assimilation as best guesses of the true state of the atmosphere. However, Analysis is created in real time as the operational best predictor of the atmosphere. ERA5 reanalysis is a dataset created in retrospect, combining observations and ECMWF’s high resolution (HRES) model predictions over twelve hour windows. In general, ERA5 data can be assumed to be closer to the truth than Analysis data, but operational models benefit from training with Analysis data. In either case, each model should be evaluated against the respective data set it was trained on, with significant differences only present for early lead times (WeatherBench, 2023). For this work, we have only considered ERA5 reanalysis data.
@@
ERA5 data exists for each hour since 1940, however all current models only train on data since 1979. This year marks a significant jump in accurate weather observation, particularly for the Southern Hemisphere, relative to previous years, as global satellite weather monitoring was introduced (Uppala, 2007). Additionally, most models consider a down-sampled version, with data every six hours at 00:00, 06:00, 12:00, and 18:00 UTC. Forecasts are initialised either on this schedule, or on a twelve hour cycle 00:00, 12:00. Predictions are then given every six or twelve hours. Forecasts are split between providing a zero-time initialisation or not, i.e. giving a prediction for its initialisation time. When comparing models with different zero-time forecasting approaches, this forecast was removed.
@@
ERA5 reanalysis data between 1979 and roughly three months behind the present exists in completion without missing data at a 0.25 latitude-longitude coordinate resolution. It can be downloaded in its entirety from the Copernicus Climate Data store (Hersbach et al., 2023). Due to the extensive work in creating this dataset, there is no need to address any data integrity issues. Predictions from WeatherBench models can be downloaded from the WeatherBench platform. Certain models do have results missing for certain variables or resolutions, along with missing predictions. Variables or resolutions with missing data were not considered. The signature kernel method can handle missing data via the time augmentation, and so no further action was taken.
Appendix I Generative model architecture of multi-step probabilistic forecaster
We employ a U-NET architecture for the generative network. U-NET consists of encoders, which are a series of convolution layers followed by max pooling. Each level increases the number of feature maps, allowing the model to learn complex features. A decoder follows, which upsamples the features along with concatenating the results of previous encoders using skip connections. This allows the decoder to reuse features from earlier layers, preserving both low-frequency and high-frequency information. Additionally, we consider a bottleneck layer between the encoder and decoder where the latent variables are introduced, to make this process generative. We consider a U-NET architecture with depths of 32, 64, 128, and 256, respectively, illustrated in I.5. We implement the same window size of 10 previous time steps, and apply the same sliding window prequential generation. Therefore, the U-NET takes a time input of size 10, and gives an output for size 1, across the 64 by 32 spatial coordinates.
Figure I.5: Diagram of the implemented U-NET model architecture. The input size is equivalent to the window size of 10. The output size is 1. The U-NET has depths of 32, 64, 128, and 256. The latent variables, introduced between the encoder and decoder, are denoted by Z.
Appendix J Additional results for training multi-step probabilistic forecaster
The following figures illustrate the performance of the signature kernel trained models across path length. Figures J.6 and J.7 showcase ensemble spread for one realisation, with random variations between members to capture the forecast distribution. Most discrepancies exist in the temperate zone where there is distinct boundary behaviour with polar regions. Remaining figures showcase one ensemble member across respective path lengths of 5, 10, and 15. Errors are minimal for each initial forecast and build with increasing path length as expected.
Figure J.6: Ensemble predictions for one initialisation generated from the path length 2 model.
Figure J.7: Ensemble differences for one initialisation generated from the path length 2 model.
Figure J.8: Path comparison for one ensemble member generated from the path length 5 model.
Figure J.9: Path differences for one ensemble member generated from the path length 5 model.
Figure J.10: Path comparison for one ensemble member generated from the path length 10 model.
Figure J.11: Path differences for one ensemble member generated from the path length 10 model.
Figure J.12: Path comparison for one ensemble member generated from the path length 15 model.
Figure J.13: Path differences for one ensemble member generated from the path length 15 model.
Appendix K Limitations
Several limitations of the current work have been collected below and are directions for future research.
As discussed in Section 4.3, numerical instability remains the one of the largest practical constraints on this method. Existing theoretical guidance on scaling and data augmentations has only been validated up to path dimensions of $d=8$, while this work considers dimensions of up to 2048 (Morrill et al., 2021). Instability has been resolved in the tested settings using an empirically founded scaling fix, and the RBF static kernel with $\sigma=1$, which may not be the most appropriate for the weather setting. Additionally, the choice of augmentation hyperparameters, such as window type and the choice not to lead lag, is informed by existing empirical evidence rather than optimised for weather data. Investigating the role of these choices would be a significant and valuable extension to this work.
The Goursat PDE solver used to compute the signature kernel requires dense signatures, meaning paths where signature terms are generally non-zero (Salvi et al., 2021). Specific weather variables, such as rainfall, are zero inflated and produce sparse signatures. This restricts the current implementation to variables with dense path behaviour, such as geopotential and temperature, but the extension to precipitation remains an open problem.
Spatial correlations are still limited by the problem of incorporating latitude weighting. The globe must be decomposed into smaller regions, either latitude slices or patches, in order to appropriately weight the signature kernel score. This necessarily discards cross-latitude or global spatial correlations. As a result, the signature kernel score can not fully exploit the global spatial structure that the method is theoretically capable of encoding.
The training scope is a further limitation. Training experiments are conducted exclusively on geopotential at 500 hPa at the coarsest WeatherBench resolution of 64 by 32, with a lightweight model. Nonetheless, with our primary aim to propose the signature kernel score as a diagnostic, we have demonstrated the reasonable success of our approach. An extension to benchmark our approach with state-of-the-art models, higher resolution data, and multi-variable forecasting would be required to draw conclusions on the method’s performance for training in real-world scenarios.
Finally, we make assumptions of stationarity and mixing conditions for the underlying process to guarantee convergence for the prequential scoring rule minimiser (Pacchiardi et al., 2024). Weather data exhibits non-stationary behaviour with seasonal cycles and longer term climate trends, which means these assumptions may only be roughly satisfied. The practical impact of this violation has not been formally evaluated in this setting.